# Verifying Well-Posedness of Linear PDEs using Convex Optimization

### 01 Apr 2026 v1

@@ Abstract
Ensuring that a PDE model is well-posed is a necessary precursor to any form of analysis, control, or numerical simulation. Although the Lumer-Phillips theorem provides necessary and sufficient conditions for well-posedness of dissipative PDEs, these conditions must hold only on the domain of the PDE – a proper subspace of $L_{2}$ – which can make them difficult to verify in practice. In this paper, we show how the Lumer–Phillips conditions for PDEs can be tested more conveniently using the equivalent Partial Integral Equation (PIE) representation. This representation introduces a fundamental state in the Hilbert space $L_{2}$ and provides a bijection between this state space and the PDE domain. Using this bijection, we reformulate the Lumer-Phillips conditions as operator inequalities on $L_{2}$. We show how these inequalities can be tested using convex optimization methods, establishing a least upper bound on the exponential growth rate of solutions. We demonstrate the effectiveness of the proposed approach by verifying well-posedness for several classical examples of parabolic and hyperbolic PDEs.
@@ I INTRODUCTION
Well-Posedness is a fundamental property of any dynamic model, ensuring existence and regularity of solutions. For ordinary differential equations, the problem is relatively simple. Any linear ordinary differential equation is well-posed and even for nonlinear ordinary differential equations, existence of a non-increasing Lyapunov function is sufficient to establish this property. Well-posedness of Partial Differential Equations (PDEs), however, is more nuanced. Specifically, uniqueness of solutions necessitates the imposition of boundary conditions (and the associated spatial Sobolev regularity required to ensure existence of such boundary values). The well-posedness condition, then, requires that the boundary conditions (BCs) are tight enough to avoid multiple solutions, while not being so tight as to preclude any solution. This process of ensuring that the BCs are compatible with the PDE is often a difficult time-consuming process, requiring careful ad hoc analysis and integration of domain (BCs) and generator (PDE). In this paper, we propose a convex optimization algorithm for testing well-posedness which applies to a large class of linear PDEs and BCs, while avoiding the need for extensive ad hoc analysis.
@@
where $\mathbf{u}(t)\in D(A)$ denotes the state of the system, defined on a Sobolev subspace $D(A)\subseteq H$ of a Hilbert space $H$, and where $A:D(A)\to H$ is a linear differential operator. Many physical phenomena can be modeled by dynamical systems of this form. Such models provide simplified descriptions of complex processes (e.g. physical, chemical, biological, etc.), enabling analysis of system properties such as stability and facilitating the synthesis of controllers and observers.
@@
A PDE model of the form in Eqn. (1) is well posed if it admits a unique solution that depends continuously on the initial state $\mathbf{x}$ [1]. A common framework for formulating well-posedness of PDEs is the theory of $C_{0}$-semigroups [2]. These semigroups generalize the matrix exponential to infinite-dimensional settings, allowing solutions of (1) to be characterized by the $C_{0}$-semigroup generated by $A$. Within this framework, the question of well-posedness is whether the operator $A$ generates a $C_{0}$-semigroup, and substantial research has been devoted to establishing generation theorems which provide conditions under which such a semigroup exists.
@@
While necessary and sufficient conditions for generation of $C_{0}$-semigroups have been established in the form of the Hille–Yosida theorem [^1] Also known as the Feller–Miyadera–Phillips theorem, for the more general formulation derived in [3, 4, 5]. [6, 7], the formulation of these conditions is abstract and verification is challenging. Consequently, most well-posedness results use the stronger conditions proposed in the Lumer–Phillips theorem [8]. These conditions are more closely related to the well-posedness test for nonlinear ODEs based on existence of non-increasing Lyapunov function. Specifically, the Lumer-Phillips theorem states that $A$ generates a contraction semigroup – implying that some norm of solutions in some Hilbert space does not increase with time – if and only if $A$ satisfies $\left\langle\mathbf{x},A\mathbf{x}\right\rangle_{H}\leq 0$ for all $\mathbf{x}\in D(A)$, and $\lambda I-A$ is surjective for some $\lambda>0$. Modifications of the Lumer–Phillips theorem have also been proposed for other classes of systems, including time-varying PDEs [9, 10] and partial differential–algebraic equations [11, 12].
@@
Despite being widely used to verify well-posedness of PDEs, application of the Lumer–Phillips theorem is typically tailored to the specific system under consideration. Development of a universal framework for testing well-posedness is complicated by the lack of a universal parameterization of operators $A$ defining PDEs on different domains $D(A)$. As a result, substantial analysis is often required to test the conditions of the Lumer–Phillips theorem. For instance, dissipativity is commonly verified using techniques such as integration by parts and the Cauchy–Schwarz inequality to establish that $\left\langle\mathbf{x},A\mathbf{x}\right\rangle_{H}\leq 0$ for all $\mathbf{x}\in D(A)$. Similarly, surjectivity of $\lambda I-A$ is frequently tested by explicitly inverting the operator $\lambda I-A$ or $A$, which becomes highly nontrivial for more complex systems.
@@
To develop a more general, computational approach to well-posedness analysis of PDEs, we propose to use the Partial Integral Equation (PIE) representation. This PIE representation provides an equivalent state-space formulation of PDEs by establishing a bijection between the PDE domain and the Hilbert space $H=L_{2}$. Under suitable rank conditions on the boundary conditions defining $D(A)$ (see [13], Sec. XIV.3, Thm. 3.1), such a bijection exists and belongs to a class of Partial Integral (PI) operators of the form
$\mathcal{T}\mathbf{v}(s):=\int_{0}^{s}\boldsymbol{T}_{1}(s,\theta)\mathbf{v}(\theta)d\theta+\int_{s}^{1}\boldsymbol{T}_{2}(s,\theta)\mathbf{v}(\theta)d\theta,$
where the kernels $\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ are polynomial and can be computed explicitly from the boundary conditions defining $D(A)$. Using this operator, $\mathcal{T}:L_{2}\to D(A)$ is bijective, yielding an equivalent representation of the PDE in (1) as
@@
where $\mathcal{A}:=A\circ\mathcal{T}^{-1}$ is also a PI operator. It has been shown that a broad class of linear ODE–PDE systems can be represented equivalently as PIEs of this form [14, 15], where $\mathbf{u}(t)$ satisfies the PDE if and only if $\mathbf{u}(t)=\mathcal{T}\mathbf{v}(t)$ for some solution $\mathbf{v}(t)$ of the PIE.
Within the PIE framework, various analysis and synthesis problems – including stability analysis as well as observer and controller design – can be formulated as convex optimization problems over PI operator variables, known as Linear PI Inequalities (LPIs). Since PI operators admit polynomial parameterizations, these LPIs can be solved using semidefinite programming methods through the PIETOOLS software suite [16].
Despite the growing use of the PIE representation for analysis and control of PDE systems, a corresponding generation theorem has not yet been formulated in the PIE framework. The presence of the operator $\mathcal{T}$ on the left-hand side of the PIE dynamics prevents direct application of classical results such as the Lumer–Phillips theorem, raising the question of how well-posedness conditions should be formulated in terms of the operators $\{\mathcal{T},\mathcal{A}\}$. Although related generation results exist for descriptor systems of the form $\frac{d}{dt}E\mathbf{x}(t)=A\mathbf{x}(t)$ [11, 12], these results still assume $A$ to be defined on a domain $D(A)$, leading to conditions that are not significantly easier to verify than the classical generator conditions. This leaves the challenge of how to exploit the bijective nature of $\mathcal{T}$ and the fact that $\mathcal{A}\in\mathcal{L}(H)$ to establish a simplified generation theorem, of which the conditions can be tested numerically.
In this paper, we address this challenge by deriving well-posedness conditions directly in the PIE representation. First, in Section III, we parameterize a class of 1D, linear PDEs admitting an equivalent PIE representation, and formulate the problem of verifying well-posedness on $L_{2}$ in terms of the associated PIE. In Section IV, we then formulate a Lumer–Phillips theorem for PIEs, obtaining necessary and sufficient conditions for the PIE to generate a quasicontraction semigroup. Finally, in Section V, we use this result to show how well-posedness of a PDE can be tested in the PIE representation by solving an LPI, further establishing a bound on the exponential growth rate of solutions. We apply this test to several classical PDE examples in Section VI, using the PIETOOLS software to verify well-posedness and compute tight bounds on the growth rate of solutions.
@@ II-A Notation
Define $\mathbb{R}_{+}:=[0,\infty)$. For $d,n\in\mathbb{N}$, denote by $L_{2}^{n}[a,b]$ the Hilbert space of $\mathbb{R}^{n}$-valued square-integrable functions on $[a,b]$, with the standard inner product $\left\langle\cdot,\cdot\right\rangle_{L_{2}}$, and where we omit the domain when clear from context. For $\mathcal{P}\in L_{2}^{n}[a,b]\to L_{2}^{n}[a,b]$ symmetric positive definite, define the weighted inner product $\left\langle\mathbf{x},\mathbf{y}\right\rangle_{\mathcal{P}}:=\left\langle\mathbf{x},\mathcal{P}\mathbf{y}\right\rangle_{L_{2}}$ for $\mathbf{x},\mathbf{y}\in L_{2}^{n}$. Define the Sobolev subspace
$W_{2}^{d,n}[0,1]:=\{\mathbf{u}\in L_{2}^{n}\mid\partial_{s}^{k}\mathbf{u}\in L_{2}^{n}[0,1],~\forall k\in\{1,\ldots,d\}\}.$
For a Hilbert space $H$, denote by $\mathcal{L}(H)$ the space of bounded linear operators on $H$ (with operator norm induced by $\left\lVert{\cdot}\right\rVert_{H}$), and let $I\in\mathcal{L}(H)$ denote the identity operator. For $\mathcal{P}\in\mathcal{L}(H)$, denote by $\textnormal{ran}(\mathcal{P})$ and $\textnormal{ker}(\mathcal{P})$ the range and nullspace of $\mathcal{P}$, respectively.
@@ II-B Strongly Continuous Semigroups
It is well-established that if $A$ is the infinitesimal generator of a $C_{0}$-semigroup, then solutions to the PDE in (1) defined by AA take the form $\mathbf{u}(t)=e^{tA}\mathbf{x}$, for initial state $\mathbf{x}\in D$ [17]. Accordingly, in the following section, well-posedness of the PDE will be defined in terms of generation of a $C_{0}$-semigroup.
@@ III Well-Posedness in the PIE Representation
The main contribution of this paper is a convex optimization program for verifying well-posedness of a class of linear PDEs. This optimization program will be established using the equivalent Partial Integral Equation (PIE) representation of the PDE, modeling the highest-order spatial derivative of the PDE state. In this section, we recall how this PIE representation may be constructed for a broad class of linear, 1D PDEs, formulating the problem of well-posedness analysis in terms of the associated PIE representation.
@@ III-A Bijection between PDE Domain and $L_{2}$
where $B_{i,j},C_{i,j}\in\mathbb{R}^{n\times n}$ are such that $B=[B_{i,j}]$ and $C=[C_{i,j}]$. We say that $D$ is PIE compatible if the matrix $K:=B+C\boldsymbol{Q}(b-a)$ is invertible, where
@@
Many common boundary conditions (e.g., Dirichlet, Robin, mixed) yield PIE compatible domains. The following result, a corollary of Thm. 3.1 in Sec. XIV.3 of [13], shows that for any such PIE compatible domain, the operator $\partial_{s}^{d}:D\to L_{2}^{n}[a,b]$ is invertible.
@@
for $\boldsymbol{Q}(s)$ and $K$ as in Defn. 2, and $E_{1}=\begin{bmatrix}I_{n}&0_{n}&\cdots&0_{n}\end{bmatrix}^{T}$ and $E_{d}=\begin{bmatrix}0_{n}&\cdots&0_{n}&I_{n}\end{bmatrix}^{T}$.
Cor. 3 shows that if a PDE domain $D\subseteq L_{2}^{n}$ is PIE compatible as in Defn. 2, then there exists a unique solution to the boundary value problem
$\partial_{s}^{n}\mathbf{u}=\mathbf{v},\qquad\mathbf{u}\in D,$ (2)
for all $\mathbf{v}\in D$. Moreover, this solution can be explicitly constructed as $\mathbf{u}=\mathcal{T}\mathbf{v}$, where $\mathcal{T}$ is a bounded, linear operator on $L_{2}^{n}$. In this manner, for a linear 1D PDE, PIE compatibility of the PDE domain already establishes well-posedness of the underlying boundary value problem in (2). This will greatly simplify well-posedness analysis of the full PDE, by using the operator $\mathcal{T}$ to establish a state space representation of the PDE in terms of $\mathbf{v}(t)\in L_{2}^{n}$—as a PIE.

### 02 Apr 2026 v2

@@ Abstract
Ensuring that a PDE model is well-posed is a necessary precursor to any form of analysis, control, or numerical simulation. Although the Lumer–Phillips theorem provides necessary and sufficient conditions for well-posedness of dissipative PDEs, these conditions must hold only on the domain of the PDE – a proper subspace of $L_{2}$ – which can make them difficult to verify in practice. In this paper, we show how the Lumer–Phillips conditions for PDEs can be tested more conveniently using the equivalent Partial Integral Equation (PIE) representation. This representation introduces a fundamental state in the Hilbert space $L_{2}$ and provides a bijection between this state space and the PDE domain. Using this bijection, we reformulate the Lumer–Phillips conditions as operator inequalities on $L_{2}$. We show how these inequalities can be tested using convex optimization methods, establishing a least upper bound on the exponential growth rate of solutions. We demonstrate the effectiveness of the proposed approach by verifying well-posedness for several classical examples of parabolic and hyperbolic PDEs.
@@ I INTRODUCTION
Well-posedness is a fundamental property of any dynamic model, ensuring existence and regularity of solutions. For Ordinary Differential Equations (ODEs), the problem is relatively simple. Any linear ODE is well-posed, and even for nonlinear ODEs, existence of a non-increasing Lyapunov function is sufficient to establish this property. Well-posedness of Partial Differential Equations (PDEs), however, is more nuanced. Specifically, uniqueness of solutions necessitates the imposition of boundary conditions (and the associated spatial Sobolev regularity required to ensure existence of such boundary values). The well-posedness condition, then, requires that the Boundary Conditions (BCs) are sufficiently restrictive to avoid multiple solutions, while not being so tight as to preclude any solution. This process of ensuring that the BCs are compatible with the PDE is often a difficult, time-consuming process, requiring careful ad hoc analysis and integration of domain (BCs) and generator (PDE). In this paper, we propose a convex optimization algorithm for testing well-posedness, which applies to a large class of linear PDEs and BCs, while avoiding the need for extensive ad hoc analysis.
@@
where $\mathbf{u}(t)\in D$ denotes the state of the system, defined on a Sobolev subspace $D\subseteq H$ of a Hilbert space $H$, and where $A:D\to H$ is a linear differential operator. Many physical phenomena can be modeled by dynamical systems of this form. Such models provide simplified descriptions of complex processes (e.g. physical, chemical, biological, etc.), enabling analysis of system properties such as stability and facilitating the synthesis of controllers and observers.
@@
A PDE model of the form in Eqn. (1) is well-posed if it admits a unique solution that depends continuously on the initial state, $\mathbf{x}$ [1]. A common framework for formulating well-posedness of PDEs is the theory of $C_{0}$-semigroups [2]. These semigroups generalize the matrix exponential to infinite-dimensional settings, allowing solutions of (1) to be characterized by the $C_{0}$-semigroup generated by $A$. Within this framework, the question of well-posedness is whether the operator $A$ generates a $C_{0}$-semigroup, and substantial research has been devoted to establishing generation theorems which provide conditions under which such a semigroup exists.
@@
While necessary and sufficient conditions for generation of $C_{0}$-semigroups have been established in the form of the Hille–Yosida theorem [^1] Also known as the Feller–Miyadera–Phillips theorem, for the more general formulation derived in [3, 4, 5]. [6, 7], the formulation of these conditions is abstract and verification is challenging. Consequently, most well-posedness results use the stronger conditions proposed in the Lumer–Phillips theorem [8]. These conditions are more closely related to the well-posedness test for nonlinear ODEs based on existence of a non-increasing Lyapunov function. Specifically, the Lumer–Phillips theorem states that $A$ generates a contraction semigroup (implying that some norm of solutions in some Hilbert space, $H$, does not increase with time) if and only if $A$ satisfies $\left\langle\mathbf{x},A\mathbf{x}\right\rangle_{H}\leq 0$ for all $\mathbf{x}\in D$, and $\lambda I-A:D\rightarrow H$ is surjective for some $\lambda>0$. Modifications of the Lumer–Phillips theorem have also been proposed for other classes of systems, including time-varying PDEs [9, 10] and partial differential–algebraic equations [11, 12].
@@
Despite being widely used to verify well-posedness of PDEs, application of the Lumer–Phillips theorem is typically tailored to the specific system under consideration. Development of a universal framework for testing well-posedness is complicated by the lack of a universal parameterization of operators $A$ defining PDEs on different domains $D$. As a result, substantial analysis is often required to test the conditions of the Lumer–Phillips theorem. For instance, dissipativity is commonly verified using techniques such as integration by parts and the Cauchy–Schwarz inequality to establish that $\left\langle\mathbf{x},A\mathbf{x}\right\rangle_{H}\leq 0$ for all $\mathbf{x}\in D$. Similarly, surjectivity of $\lambda I-A$ is frequently tested by explicitly inverting the operator $\lambda I-A$.
@@
To develop a universal computational approach to well-posedness analysis of PDEs, we consider the Partial Integral Equation (PIE) representation. This PIE representation provides an equivalent state-space formulation of PDEs by establishing a bijection between the PDE domain and the Hilbert space $L_{2}$. Under suitable rank conditions on the BCs defining $D$ (see [13], Sec. XIV.3, Thm. 3.1), such a bijection exists and has the form
$(\mathcal{T}\mathbf{v})(s):=\int_{0}^{s}\boldsymbol{T}_{1}(s,\theta)\mathbf{v}(\theta)\,d\theta+\int_{s}^{1}\boldsymbol{T}_{2}(s,\theta)\mathbf{v}(\theta)\,d\theta,$
where the kernels $\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ are polynomial and can be computed explicitly from the BCs defining $D$. We refer to operators of this type as Partial Integral (PI) operators [^2] PI operators are formally defined in Subsec. V-B and may be augmented with multiplier operators. In both cases (with and without multipliers), the set of PI operators form a *-algebra under composition, with analytic formulae for algebraic operations expressed using the associated kernels and multipliers.. Using this bijection, $\mathcal{T}:L_{2}\to D$, an equivalent representation of the PDE in (1) can be expressed as a PIE,
@@
where $\mathcal{A}:=A\circ\mathcal{T}$ is also a PI operator. It has been shown that a broad class of linear ODE–PDE systems can be represented equivalently as PIEs of this form [14, 15], where $\mathbf{u}(t)$ satisfies the PDE if and only if $\mathbf{u}(t)=\mathcal{T}\mathbf{v}(t)$ for some solution $\mathbf{v}(t)$ of the PIE.
Within the PIE framework, various analysis and synthesis problems – including stability analysis as well as observer and controller design – can be formulated as convex optimization problems over PI operator variables. Such optimization problems are termed Linear PI Inequalities (LPIs). Since PI operators may be parameterized by their polynomial kernels and multipliers, these LPIs can be solved using semidefinite programming methods through the PIETOOLS software suite [16].
Despite the growing use of the PIE representation for analysis and control of PDE systems, a corresponding generation theorem has not yet been formulated in the PIE framework. The challenge is that the operator $\mathcal{T}$ on the left-hand side of the PIE dynamics prevents direct application of classical results such as the Lumer–Phillips theorem. In particular, while the condition $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ can be readily tested in the PIE representation as an LPI, $\mathcal{T}^{*}\mathcal{A}+\mathcal{A}^{*}\mathcal{T}\preceq 0$, an analogous test for surjectivity of $\lambda I-A$ for $\lambda>0$ has not been developed. Although generation results do exist for descriptor systems, $\frac{d}{dt}E\mathbf{x}(t)=A\mathbf{x}(t)$ [11, 12], these results still assume $A$ to be defined on a domain $D$, leading to conditions that remain difficult to verify. This raises the challenge of using the bijective nature of $\mathcal{T}$ and algebraic properties of $\{\mathcal{T},\mathcal{A}\}$ to formulate a simplified generation theorem, of which the conditions can be tested numerically.
In this paper, we address this challenge by deriving well-posedness conditions directly in the PIE representation and expressing them as LPIs. First, in Section III, we parameterize a class of 1D linear PDEs admitting an equivalent PIE representation, and pose well-posedness on $L_{2}$ in terms of the PIE. In Section IV, we then establish a Lumer–Phillips theorem for PIEs, yielding necessary and sufficient conditions for generation of a contraction semigroup. Finally, in Section V, we show how these conditions can be tested via an LPI, further establishing a bound on the exponential growth rate of solutions. We apply this test to several classical PDE examples in Section VI, using the PIETOOLS software to verify well-posedness and compute tight bounds on the growth rate of solutions.
@@ II-A Notation
$\mathbb{R}[x]^{n\times m}$ is the space of matrix-valued polynomials on $x$. For $d,n\in\mathbb{N}$, denote by $L_{2}^{n}[a,b]$ the Hilbert space of $\mathbb{R}^{n}$-valued square-integrable functions on $[a,b]$, with the standard inner product $\left\langle\cdot,\cdot\right\rangle_{L_{2}}$, and where we omit the domain when clear from context. Define the Sobolev subspace
$W_{2}^{d,n}[a,b]:=\{\mathbf{u}\in L_{2}^{n}\mid\partial_{s}^{k}\mathbf{u}\in L_{2}^{n}[a,b],~\forall k\in\{1,\ldots,d\}\}.$
For a Hilbert space $H$, denote by $\mathcal{L}(H)$ the space of bounded linear operators on $H$ (with operator norm induced by $\left\langle\cdot,\cdot\right\rangle_{H}$), and let $I\in\mathcal{L}(H)$ denote the identity operator. For coercive $\mathcal{P}\in\mathcal{L}(L_{2}^{n}[a,b])$, so that $\mathcal{P}\succeq\epsilon I$ for some $\epsilon>0$, define the Hilbert space $L_{2,\mathcal{P}}^{n}[a,b]$ to be $L_{2}^{n}[a,b]$, equipped with the weighted inner product $\left\langle\mathbf{x},\mathbf{y}\right\rangle_{L_{2,\mathcal{P}}}:=\left\langle\mathbf{x},\mathcal{P}\mathbf{y}\right\rangle_{L_{2}}$. For $\mathcal{P}\in\mathcal{L}(H)$, denote by $\textnormal{ran}(\mathcal{P})$ and $\textnormal{ker}(\mathcal{P})$ the range and nullspace of $\mathcal{P}$, respectively.
@@ III Well-Posedness in the PIE Representation
The main contribution of this paper is a convex optimization program for verifying well-posedness of a class of linear PDEs. This optimization program will be established using the equivalent Partial Integral Equation (PIE) representation of the PDE, where if $D\subseteq W^{d,n}$, the PIE state is the $d^{th}$ order spatial derivative of the PDE state. In this section, we recall how this PIE representation may be constructed for a broad class of linear 1D PDEs. We then formulate the problem of well-posedness analysis in terms of the associated PIE representation.
@@ III-A Bijection between PDE Domain and $L_{2}$
where $B_{i,j},C_{i,j}\in\mathbb{R}^{n\times n}$ are concatenated as $B=[B_{i,j}]$ and $C=[C_{i,j}]$. We say that $D$ is PIE compatible if the matrix $K:=B+C\boldsymbol{Q}(b-a)$ is invertible, where
@@
Most common boundary conditions (e.g., Dirichlet, Robin, mixed) yield PIE compatible domains [^3] The case of periodic boundary conditions is not prima facie PIE compatible, but can be formulated in the PIE representation as in [17].. The following result, a corollary of Thm. 3.1 in Sec. XIV.3 of [13], shows that for any such PIE compatible domain, the operator $\partial_{s}^{d}:D\to L_{2}^{n}[a,b]$ is invertible.
@@
for $\boldsymbol{Q}(s)$, $K$, and $C$ as in Defn. 2, and
$E_{1}:=\begin{bmatrix}I_{n}&0_{n}&\!\cdots\!&0_{n}\end{bmatrix}^{T},\quad E_{d}:=\begin{bmatrix}0_{n}&\!\cdots\!&0_{n}&I_{n}\end{bmatrix}^{T}.$
Cor. 3 shows that if a PDE domain $D\subseteq W_{2}^{d,n}$ is PIE compatible as in Defn. 2, then for any $\mathbf{v}\in L_{2}^{n}$, there exists a unique solution, $\mathbf{u}$, to the boundary value problem
$\partial_{s}^{d}\mathbf{u}=\mathbf{v},\qquad\mathbf{u}\in D.$ (2)
Moreover, this solution can be explicitly constructed as $\mathbf{u}=\mathcal{T}\mathbf{v}$, where $\mathcal{T}$ is a bounded, linear operator on $L_{2}^{n}$. In this manner, for a linear 1D PDE, PIE compatibility of the PDE domain already establishes well-posedness of the underlying boundary value problem in (2). This will greatly simplify well-posedness analysis of the full PDE, by using the operator $\mathcal{T}$ to establish a PIE state space representation of the PDE in terms of PIE state $\mathbf{v}(t)\in L_{2}^{n}$.